# Adaptive Designs in Fast-Track Registration Processes

### v1

@@ 4.2 Adaptive Study Designs for Fast-Track Processes
Similar to Subsection 4.1, for achieving an overall power of at least $1-\beta$, the information $I_{1}$ must be strictly greater than $I_{1,\min}$. Furthermore, it can be shown that for $\delta>\delta_{\text{rel}}$ and $\eta_{f}$ and all $I_{1}>I_{1,\min}$ with $\mathcal{S}_{A}(0,I_{1})\leq 1-\beta$ there exists a unique minimum information $I_{2,\min}$ such that $\mathcal{S}_{A}(I_{2,\min},I_{1})=1-\beta$. For $I_{1}$ with $\mathcal{S}_{A}(0,I_{1})\geq 1-\beta$ there is no minimum information needed for the second stage. In the latter case, we set $I_{2,\min}=0$. For proof we refer to the Appendix A.
Example (continued).
For our example fast-track procedure, the calculation of the information in (12) means that for an observed healthcare effect $\hat{\vartheta}_{1}$ such that the corresponding noncentrality parameter fulfils $z_{1}=\sqrt{I_{1}}\hat{\vartheta}_{1}\geq z_{f}$ the second stage sample size per group is given by (with $n_{2,\min}:=2\hat{\sigma}^{2}I_{2,\min}$)
$n_{2}^{A}(z_{1}):=2\hat{\sigma}^{2}I_{2}^{A}(z_{1})=\max\left\{n_{2,\min},2\hat{\sigma}^{2}\{\Phi^{-1}(1-\beta)+\Phi^{-1}(1-A(z_{1}))\}^{2}/\hat{\vartheta}_{1}^{2}\right\}.$
The black dashed and blue solid lines in Figure 3 show the minimum required sample sizes relative to $I_{\delta}$, i.e. $t_{\xi}(I_{2,\min})$, for the two different $\xi$ with the conditional error functions $A_{I,z_{f}}$ and $A_{F,z_{f}}$, respectively. It can be seen that, with regard to the minimum needed information for the second stage, the control of the overall power ($0.81-\beta=$ 0.8) at $\delta=\xi\delta_{\text{rel}}$ comes for the adaptive designs with a much smaller requirement for the minimum information than for the non-adaptive design. For the adaptive designs, the minimum second stage information becomes quickly less than the information $I_{\delta}$ required for the fixed-sample size test and is less than 50% of this information if the interim information is $0.55\cdot I_{\delta}$ or more. For the example fast-track procedure recall that for $\xi=2$ the relative information $t_{\xi}(I_{1})=n_{1}/n_{\delta}=63/105=0.6$ corresponds to the chosen first stage sample size per group $n_{1}=63$. According to Figure 3(b) the minimum relative information needed for the second stage is now $t_{\xi}(I_{2,\min})=0.33$ and $t_{\xi}(I_{2,\min})=0.29$ for the inverse normal method and Fisher’s product test, respectively. This implies $n_{2,\min}=0.33\cdot n_{\delta}\approx 35$ and $n_{2,\min}=0.29\cdot n_{\delta}\approx 31$, respectively. These sample sizes are much lower than $n_{2,\min}=148$ needed for the non-adaptive design.
4.3 Comparison of different study designs
We now want to compare the efficiency of conducting an adaptive design (with conditional error function $A\leq 0.5$) to a design with two separate stages with regard to the expected and maximum information under the same assumptions as in the previous sections. In particular, this implies that for a given interim information $I_{1}$ the applicant achieves the overall power of $1-\beta$ at the a priori assumed effect $\delta=\xi\cdot\delta_{\text{rel}}$ by choice of the required minimum second stage information $I_{2,\min}$. We will compare the expected second stage information in dependence of $I_{1}>I_{1,\min}$ assuming that the true effect is $\delta$. Additionally, we will compare the maximum information for the second stage. Like for the first stage, expected and maximum informations will be compared for different $\xi$ and again relative to $I_{\delta}$, i.e. on the $t_{\xi}$-scale for a most general comparison which only depends on $\alpha_{c}$, $\alpha$, $\beta$ and $\xi$.
Figure 4 shows for $\xi=1.75$ and $\xi=2$ the (relative) mean information for the second stage for the two adaptive designs with conditional error functions $A_{I,z_{f}}$ and $A_{F,z_{f}}$ and the non-adaptive design. It can be seen that the adaptive designs have for all $t_{\xi}(I_{1})$ with $t_{\xi}(I_{1,\min})<t_{\xi}(I_{1})\leq t_{\xi}(I_{1,\max})$ a lower (relative) mean information. Furthermore, for the adaptive designs for the most $t_{\xi}(I_{1})$, the relative mean information is (clearly) less than $1$. Thus, the mean information needed for the second stage is less than the information needed for a (single) fixed-size sample test with power $1-\beta$ (direct application for permanent registration). Furthermore, it can be numerically calculated that the mean information over both stages is not much larger than $1$. The corresponding Figures can be found in Appendix C.1.
(a) $\xi=1.75$.
(b) $\xi=2$.
Figure 4: Visualisation of the mean information for the second stage for adaptive and non-adaptive designs.
The maximum informations of the considered designs are shown in Figure 5. Figure 4 and Figure 5 have in common that if $t_{\xi}(I_{1})$ comes closer to $t_{\xi}(I_{1,\min})$, the mean and maximum information for the second stage become large and tend to infinity. This is because the minimum information $t_{\xi}(I_{2,\min})$ tends to infinity when $t_{\xi}(I_{1})$ comes closer to the critical value $t_{\xi}(I_{1,\min})$ as described in the previous subsections. As it can be seen from Figure 5(b) for $\xi=2$, the maximum information has a linear increasing part between $0.4498$ and $0.5474$ (adaptive designs) or between $0.4511$ and $0.5474$ (non-adaptive design). This is because, for $\xi=2$ for all $t_{\xi}(I_{1})\leq 0.5474$, the border $z_{f}$ is the same: It holds $z_{f}=1-\Phi^{-1}(1-\alpha_{c})=1.0364$ (i.e. the demand $p_{1}\leq\alpha_{c}$ is stricter than having a numeric relevant estimation). This implies that the maximum information is a locally linear function of $t_{\xi}(I_{1})$. For $\xi=1.75$, there is no such interval on which $z_{f}$ is constant. Thus, there is no increasing part in the maximum information in Figure 5(a). For both adaptive designs for $\xi=1.75$ and $\xi=2$ the relative maximum information remain constant for $t_{\xi}(I_{1})>1.06$ and $t_{\xi}(I_{1})>1.38$, respectively, because the conditional error functions are then constantly equal to $0.5$. Figure 5 shows that the adaptive designs are also convincing with regard to the maximum information when compared with the non-adaptive design. We analyse this result.
(a) $\xi=1.75$.
(b) $\xi=2$.
Figure 5: Visualisation of the maximum information for the second stage for adaptive and non-adaptive designs.
Recall that the maximum sample size is achieved for $z_{1}=z_{f}$ or equals $I_{2,\min}$. It seems obvious that an adaptive design with conditional error function $A$ has lower maximum information than the non-adaptive design if $A(z_{f})>\alpha$ is fulfilled. Nevertheless, since the maximum information can also be equal to $I_{2,\min}$, a formal proof is needed (see Appendix A). For our considered conditional error functions for $\xi=1.75$ and $\xi=2$ for all $I_{1}>I_{1,\min}$ it holds indeed $A_{I,z_{f}}(z_{f})>\alpha$ and $A_{F,z_{f}}(z_{f})>\alpha$ (proven in Appendix A). This gives a formal explanation for the lower maximum information of the adaptive designs in Figure 5. It is important to note that the binding futility bound enables the possibility to construct conditional error functions $A$ which fulfil for all $z_{1}\geq z_{f}$ that $A(z_{1})>\alpha$ and thus $A(z_{f})>\alpha$ while having the Type I error controlled at level $\alpha$. In the most extreme case, the conditional error function, which is constant equals ${\alpha}/{\alpha_{f}}>\alpha$, can be chosen with the binding futility bound $Z_{1}<z_{f}$. This design would probably only be accepted if the authorities were informed about it before the start of data collection of the first stage.
Example (continued).
For our example fast-track procedure for the a priori assumed $\xi=2$ and the considered first stage sample size of $n_{1}=63$ (i.e. $t_{\xi}(I_{1})=0.6$) Figure 5(b) yields $t_{\xi}(I_{2,\max})=4,2.78,3$ for the non-adaptive design, the inverse normal method and Fishers’ product test, respectively. Using the relation $t_{\xi}(I_{2,\max})=n_{2,\max}/n_{\delta}$ and the results from the previous sections we have that the sample size per group for the second stage lies between: $n_{2,\min}=148$ and $n_{2,\max}=420$ (non-adaptive design), $n_{2,\min}=35$ and $n_{2,\max}=292$ (inverse normal method), $n_{2,\min}=31$ and $n_{2,\max}=315$ (Fisher’s product test). Figure 13(b) yields $t_{\xi}(E(I_{2}))=1.41$ (non-adaptive design) and $t_{\xi}(E(I_{2}))=0.47$ (adaptive designs). This implies a second stage mean sample size per group of $E(n_{2})=1.41\cdot n_{\delta}=149$ (non-adaptive design) and $E(n_{2})=0.47\cdot n_{\delta}\approx 50$ (adaptive designs). The average sample sizes of the second stage are close to the minimum sample sizes, which implies that the distribution of the sample size is left-skewed in the sense that sample sizes close to the maximum ones are rare.
5 Combination of conditional and permanent registration
Recall from Section 3.2 that, under the assumption that the applicants cannot apply for a permanent registration after an unsuccessful attempt to obtain conditional registration, the fast-track procedure does not make sense for all $\xi\leq\xi_{\min}=1.43$. Under these assumptions, the probability of overall success, which we understand as the probability of achieving permanent registration, is given by $\mathbb{P}_{\delta}(\text{reject\leavevmode\nobreak\ }H_{0})=\mathbb{P}_{\delta}(Z_{1}\geq z_{f},\text{reject\leavevmode\nobreak\ }H_{0})$, and we have seen that for $\xi\leq\xi_{\min}=1.43$ there is no first-stage information $I_{1}$ for which an overall success probability of $0.81-\beta=$ 0.8 can be achieved, while the requirements for a conditional registration are less strict than those for a direct permanent registration (i.e., $I_{1,\min}>I_{1,\max}$). Motivated by this and to deal with an a priori assumed (or true) effect of $\delta/\delta_{\text{rel}}\leq\xi_{\min}$, we relax the assumption that a conditional registration is a requirement for permanent registration. We assume that, based on the pilot data, we can decide to either apply for conditional registration or waive this application and continue directly with the second stage for a permanent registration. We will work again with $\alpha_{c}=0.15$.
5.1 Suggestion for a combination strategy
A reasonable strategy for this is to apply for conditional registration whenever $Z_{1}\geq z_{f}$ (since the requirements are met), and otherwise to waive this application and continue the two-step approach with the second stage for permanent registration. This process is illustrated in Figure 6. In order to meet the overall type I error rate $\alpha$, we utilize adaptive design methodology and use for both cases an overall conditional error function $A$ satisfying (9). We will discuss possible choices for $A$ in the following.
[…]
@@ 5.1 Suggestion for a combination strategy
Figure 6: Combination strategy with a data driven decision to either apply or waive the application for conditional registration where the overall type I and II errors are under control.
@@
Moreover, if $Z_{1}\geq z_{f}$ (and conditional registration is achieved), we choose, as in Section 4, the second-stage sample size such that the conditional power $1-\beta$ is reached at the estimated effect estimate. As in Section 4, condition (14) can be satisfied with a sufficiently large minimum second stage information $I_{2,\min}$. In contrast to the situation in Section 4, we find now for all $\xi>1$ such a minimal second stage information. This is because, for all $\delta>0$, the left-hand side of (14) converges to 1 with increasing second-stage information.
Lemma 1 in Appendix A shows for a design with conditional error function $A$ for all $\xi\geq 1$ and $I_{1}>0$, that there exists a unique minimal information $I_{2,\min}$ such that (14) can be met with equality if $\mathcal{S}_{A}(I_{1},0)\leq(1-\beta)\cdot\mathbb{P}_{\delta}(Z_{1}\geq z_{f})$ (where $\mathcal{S}_{A}$ is as in (4.2), and if $\mathcal{S}_{A}(I_{1},0)>(1-\beta)\cdot\mathbb{P}_{\delta}(Z_{1}\geq z_{f})$ condition (14) is achieved with an arbitrary small information for the second stage. In this case, we can set $I_{2,\min}=0$.
We note that the use of the constant conditional error function, $A(z_{1})=\alpha$ for all $z_{1}$, corresponds to the standard fast-track process if $Z_{1}\geq z_{f}$. For $Z_{1}<z_{f}$ it results in a new (second-stage) study for permanent registration with information $I_{2,\text{const}}=I_{\delta}$ and without using the pilot data. A potentially more efficient approach that utilizes the pilot data also for the permanent registration (in both cases) is to use a non-constant conditional error function like $A_{I,-\infty}$ or $A_{F,-\infty}$. Note that these conditional error functions should not include the binding futility stopping rule $Z_{1}<z_{f}$, because we are testing now $H_{0}$ also if $Z_{1}<z_{f}$.
@@
$Z=\sqrt{I_{1}/(I_{1}+I_{2,\text{const}})}Z_{1}+\sqrt{I_{2,\text{const}}/(I_{1}+I_{2,\text{const}})}Z_{2}\geq\Phi^{-1}(1-\alpha)\leavevmode\nobreak\ .$ (15)
For our example fast-track procedure it holds $I_{1}=n_{1}/(2\sigma^{2})$ and $I_{2,\text{const}}=n_{2,\mathrm{const}}/(2\sigma^{2})$ with sample sizes per group $n_{1}$ and $n_{2,\mathrm{const}}$. Thus, (15) is equivalent to
$Z=\sqrt{n_{1}/(n_{1}+n_{2,\mathrm{const}})}Z_{1}+\sqrt{n_{2,\mathrm{const}}/(n_{1}+n_{2,\mathrm{const}})}Z_{2}\geq\Phi^{-1}(1-\alpha)\leavevmode\nobreak\ .$
@@
$\tilde{A}_{Z,\alpha}(z_{1}):=1-\Phi\left(\frac{\Phi^{-1}(1-\alpha)-\sqrt{I_{1}/(I_{1}+I_{2,\text{const}})}z_{1}}{\sqrt{I_{2,\text{const}}/(I_{1}+I_{2,\text{const}})}}\right),\quad z_{1}\in\mathbb{R}$
which satisfies the level condition (9). To complete the design, for the upper branch, a conditional error function must also be set such that the type I error rate does not exceed $\alpha$. Here we choose $\tilde{A}_{Z,\alpha^{\prime}}$ where $\alpha\leq\alpha^{\prime}\leq 1$ is the largest possible value such that the Type I error rate of
$A_{Z}(z_{1}):=0.5\wedge\begin{cases}\tilde{A}_{Z,\alpha^{\prime}}(z_{1}),&\text{if $z_{1}\geq z_{f}$}\\ \tilde{A}_{Z,\alpha}(z_{1}),&\text{if $z_{1}<z_{f}$}\leavevmode\nobreak\ .\end{cases}$ (16)
is equal to or less than $\alpha$ (i.e. (9) holds true). For all examples considered in the following, numerical calculations show that the significance level $\alpha$ is fully used, i.e. for $A_{Z}$ equality holds in (9). Note that the proposed strategy is based on the Müller and Schäfer method [27] with the difference that the conditional error function $A_{Z}$ is truncated here by $0.5$. In line with the proposed strategy in Figure 6, we assume that $I_{2,\min}$ and $I_{2,\text{const}}$ are chosen such that the aimed conditional success probabilities specified in Figure 6 are achieved.
In the following, we investigate and compare the performance of the conditional error functions $A_{Z}$, $A_{I,-\infty}$ and $A_{F,-\infty}$ and the non-adaptive design.
(a) Non-adaptive design.
(b) Inverse normal method.
(c) Fisher’s product test.
(d) $A_{Z}$.
@@ 5.2 Investigation of the second stage information $I_{2,\text{const}}$ for the case $Z_{1}<z_{f}$
The red dotted lines in the plots (a) to (d) of Figure 7 show for the case $Z_{1}<z_{f}$ the constant second-stage information $t_{\xi}(I_{2,\text{const}})$ in dependence of $t_{\xi}(I_{1})$ when $\xi=1.25$ and $\alpha_{c}=0.15$. The plots are for the non-adaptive design ($A\equiv\alpha$) and the adaptive designs with conditional error functions $A_{I,-\infty}$, $A_{F,-\infty}$ and the specific inverse normal conditional error function $A_{Z}$ that is line with $I_{2,\text{const}}$ defined in (16), respectively. Figure 8(a) allows a direct comparison of $t_{\xi}(I_{2,\text{const}})$ between the designs (red dotted line for non-adaptive design, black dashed line for inverse normal method, blue solid line for Fishers product test and purple long-dashed line for $A_{Z}$). For the non-adaptive design we get $I_{2,\text{const}}=I_{\delta}$ for all $t_{\xi}(I_{1})$. We can see that for the adaptive designs, the information $I_{2,\text{const}}$ decreases with $t_{\xi}(I_{1})$ and becomes smaller than $I_{\delta}$ if $t_{\xi}(I_{1})>0.5$ (for $A_{I,-\infty}$ and $A_{F,-\infty}$) or $t_{\xi}(I_{1})>0.33$ (for $A_{Z}$).
For the numerical determination of $I_{2,\text{const}}$, we used a bisection search. This requires that the conditional success probability is non-decreasing in $I_{2,\text{const}}$. For the non-adaptive designs and for the adaptive designs with $A_{I,-\infty}$ and $A_{F,-\infty}$, this can easily be seen. For $A_{Z}$, the monotonicity is proven in Lemma 3 in the appendix.
@@
For our example fast-track procedure, from now on we work with another value for the minimal clinically relevant effect, namely $\delta_{\text{rel}}=1.4$. This is also a reasonable assumption according to [17]. We further assume hypothetically that a fast-track procedure is now planned based on the results [28], which were used for the planning of the confirmatory study of the DiGA elevida. Based on [28], a reasonable and conservative a priori assumption for the healthcare effect is $\delta=1.75=1.25\cdot\delta_{\text{rel}}$, i.e. $\xi=1.25$. Additionally we continue to assume that a parallel group design for both conditional and permanent registration each with balanced samples, is used. Now, a information of $I_{\delta}:=\eta_{f}^{2}/\delta^{2}=2.8^{2}/1.75^{2}=2.56$, i.e. a per group sample size of $n_{\delta}:=2\hat{\sigma}^{2}I_{\delta}\approx 137$, is required for a fixed-size sample test with significance level $\alpha=0.025$ and power $0.81-\beta=$ 0.8. This corresponds to the per group sample size of the study [17] conducted for permanent registration of the diga elevida. If now $0.5=t_{\xi}(I_{1})=n_{1}/n_{\delta}$, i.e. $n_{1}=0.5\cdot 137=69$ would be the chosen first stage sample size per group, Figure 8(a) yields the sample size for the final stage without conditional registration $n_{2,\mathrm{const}}=t_{\xi}(I_{2,\mathrm{const}})\cdot n_{\delta}=1\cdot n_{\delta}=137$ with the non-adaptive design, the inverse normal method and Fisher’s product test, or $n_{2,\mathrm{const}}=0.9\cdot n_{\delta}\approx 124$ with $A_{Z}$.

### v3

@@ 5.1 Suggestion for a combination strategy
Figure 6: Combination strategy with a data-driven decision to either apply or abstain from the application for conditional registration, where the overall type I and II errors are under control.
@@
Moreover, if $Z_{1}\geq z_{f}$, we choose, as in Section 4, the second-stage sample size such that the conditional power $1-\beta$ is reached at the estimated effect estimate. As in Section 4, condition (15) can be satisfied with a sufficiently large minimum second stage information $I_{2,\min}$. In contrast to the situation in Section 4, we find now for all $\xi>1$ such minimal second-stage information. This is because, for all $\delta>0$, the left-hand side of (15) converges to 1 with increasing second-stage information.
In the Appendix, it is shown that, for a design with conditional error function $A$ for all $\xi\geq 1$ and $I_{1}>0$, there exists a unique minimal information $I_{2,\min}$ such that (15) can be met with equality if $\mathcal{S}_{A}(I_{1},0)\leq(1-\beta)\cdot\mathbb{P}_{\delta}(Z_{1}\geq z_{f})$, where $\mathcal{S}_{A}$ is as in (4.2). If $\mathcal{S}_{A}(I_{1},0)>(1-\beta)\cdot\mathbb{P}_{\delta}(Z_{1}\geq z_{f})$, condition (15) is achieved with an arbitrarily small information for the second stage. In this case, we can set $I_{2,\min}=0$.
We note that the use of the constant conditional error function, $A(z_{1})=\alpha$ for all $z_{1}$, corresponds to the standard fast-track process if $Z_{1}\geq z_{f}$. For $Z_{1}<z_{f}$, it results in a new (second-stage) study for permanent registration with information $I_{2,\text{const}}=I_{\delta}$ and without using the pilot data. A potentially more efficient approach that utilizes the pilot data also for the permanent registration (in both cases) is to use a non-constant conditional error function like $A_{I,-\infty}$ or $A_{F,-\infty}$. Note that these conditional error functions should not include the binding futility stopping rule $Z_{1}<z_{f}$, because we are testing now $H_{0}$ also if $Z_{1}<z_{f}$.
@@
$Z=\sqrt{I_{1}/(I_{1}+I_{2,\text{const}})}Z_{1}+\sqrt{I_{2,\text{const}}/(I_{1}+I_{2,\text{const}})}Z_{2}\geq\Phi^{-1}(1-\alpha)\penalty 10000\ .$ (16)
@@
$\tilde{A}_{Z,\alpha}(z_{1}):=1-\Phi\left(\frac{\Phi^{-1}(1-\alpha)-\sqrt{I_{1}/(I_{1}+I_{2,\text{const}})}z_{1}}{\sqrt{I_{2,\text{const}}/(I_{1}+I_{2,\text{const}})}}\right),\quad z_{1}\in\mathbb{R},$
which satisfies level condition (10). To complete the design, for the upper branch, the conditional error function must be extended in such a way that the type I error rate does not exceed $\alpha$. Here we choose $\tilde{A}_{Z,\alpha^{\prime}}$ where $\alpha\leq\alpha^{\prime}\leq 1$ is the largest possible value such that the Type I error rate of
$A_{Z}(z_{1}):=0.5\wedge\begin{cases}\tilde{A}_{Z,\alpha^{\prime}}(z_{1}),&\text{if $z_{1}\geq z_{f}$}\\ \tilde{A}_{Z,\alpha}(z_{1}),&\text{if $z_{1}<z_{f}$}\penalty 10000\ .\end{cases}$ (17)
is equal to or less than $\alpha$ (i.e. (10) holds true). For all examples considered in the following, it can be seen that the significance level $\alpha$ is fully used, i.e. for $A_{Z}$ equality holds in (10). Note that the proposed strategy is based on the Müller and Schäfer method [33] with the difference that the conditional error function $A_{Z}$ is truncated here by $0.5$. In line with the proposed strategy in Figure 6, we assume that $I_{2,\min}$ and $I_{2,\text{const}}$ are chosen such that the aimed conditional success probabilities specified in Figure 6 are achieved.
In the following, we investigate and compare the performance of the conditional error functions $A_{Z}$, $A_{I,-\infty}$, and $A_{F,-\infty}$ and the non-adaptive design.
@@ 5.2 Investigation of the second stage information $I_{2,\text{const}}$ for the case $Z_{1}<z_{f}$
The dotted lines in the plots (a) to (d) of Figure 7 show for the case $Z_{1}<z_{f}$ the constant second-stage information $t_{\xi}(I_{2,\text{const}})$ in dependence of $t_{\xi}(I_{1})$ when $\xi=1.25$ and $\alpha_{c}=0.15$. The plots are for the non-adaptive design ($A\equiv\alpha$) and the adaptive designs with conditional error functions $A_{I,-\infty}$, $A_{F,-\infty}$ and the specific inverse normal conditional error function $A_{Z}$ that is line with $I_{2,\text{const}}$ defined in (17), respectively. In the Supplement, a figure is provided that allows a direct comparison of $t_{\xi}(I_{2,\text{const}})$ between the designs. For the non-adaptive design we get $I_{2,\text{const}}=I_{\delta}$ for all $t_{\xi}(I_{1})$. We can see that for the adaptive designs, the information $I_{2,\text{const}}$ decreases with $t_{\xi}(I_{1})$ and becomes smaller than $I_{\delta}$ if $t_{\xi}(I_{1})>0.5$ (for $A_{I,-\infty}$ and $A_{F,-\infty}$) or $t_{\xi}(I_{1})>0.33$ (for $A_{Z}$).
For the numerical determination of $I_{2,\text{const}}$, we used a bisection search. This requires that the conditional success probability is non-decreasing in $I_{2,\text{const}}$. For the non-adaptive designs and for the adaptive designs with $A_{I,-\infty}$ and $A_{F,-\infty}$, this can easily be seen. For $A_{Z}$, the monotonicity is proven in the Appendix.
@@
For our example fast-track procedure, we will work with $\delta_{\text{rel}}=1.4$ from now on, which is also reasonable based on Pöttgen et al [23]. We further assume that the planning is now based on the results of Van Kessel et al [34], which were used for the planning of the confirmatory study of the DiGA elevida. Then, a reasonable and conservative a priori assumption for the healthcare effect is $\delta=1.75=1.25\cdot\delta_{\text{rel}}$, i.e. $\xi=1.25$. Additionally, we continue to assume that a parallel group design for both conditional and permanent registration, each with balanced samples, is used. Now, a information of $I_{\delta}:=\eta_{f}^{2}/\delta^{2}=2.8^{2}/1.75^{2}=2.56$, i.e. a per group sample size of $n_{\delta}:=2\hat{\sigma}^{2}I_{\delta}\approx 137$, is required for a fixed-size sample test with significance level $\alpha=0.025$ and power $1-\beta=0.8$. This corresponds to the per group sample size of the study conducted for permanent registration of the DiGA elevida [23]. If now $0.5=t_{\xi}(I_{1})=n_{1}/n_{\delta}$, i.e. $n_{1}=0.5\cdot 137=69$ would be the chosen first stage sample size per group, Figure 7 yields the sample size for the final stage without conditional registration $n_{2,\mathrm{const}}=t_{\xi}(I_{2,\mathrm{const}})\cdot n_{\delta}=1\cdot n_{\delta}=137$ with the non-adaptive design, the inverse normal method and Fisher’s product test, or $n_{2,\mathrm{const}}=0.9\cdot n_{\delta}\approx 124$ with $A_{Z}$.